Dear Richard dear all,

please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
ADS: P. Lazzarini
LBTO: G. Brusa

Total # of person: 9

W unit

1) The IRTC has been installed on the AGW unit by the LBTO crew. Cooling, pipes and electronic cabling has been intalled. A functionality test of the IRTC has been performed and a few images have been acquired showing no issues.
2) Flexure test of the W unit has been performed for different angles of the RFBG rotator in a range of about 180deg. Data have been acquired and data reduction is ongoing.
3) Reset of the power control board. The power control board of the W unit electronic, located in the CCD39 electronic box, expirienced two reset during the flexure test. The two reset happened in a similar angle range the rotator. The most probable cause of this failure is a loosen power cable (110V) connecting the CCD39 and CCD47 electronic boxes. The simplest possible fix of the problem require to unmount the front cooling plates of the boxes and check the power cable connection. This require sometime to access the electronic boxes while the AGW is mounted on the telescope, so the opportunity to do this fixing during this run is under discussion.

AO control SW

1) Development of the software interface for the controlling the alinement led
2) Development of the software for the high order off load during seeing limited and AO operations.
3) Refurbishment of the telemetry software analysis for the LBT672

LBT672a

1) the retroreflecting optics has been installed on the retro-reflector holder in the correct focal position considering the as-built dimensions of the holder, optical barrel and TS3 RoC
2) the sonic anemometer has been power cycled and it is currently working again. The wind values are read from the Data Dictionary Viewer GUI

Note:
P.Stefanini helped P.Stefanini in fixing the M3 unit. The work has been successfully completed by 2.00pm


Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback